### v2

# Pelican Soup Framework: A Theoretical Framework for Language Model Capabilities

Abstract
In this work, we aim to better understand how pretraining allows LLMs to (1) generalize to unseen instructions and (2) perform in-context learning, even when the verbalizers are irrelevant to the task. To this end, we propose a simple theoretical framework, Pelican Soup, based on the logical consistency of the training data, the notion of “reference-meaning association”, and a simple formalism for natural language processing tasks. Our framework demonstrates how linguistic and psychology studies can inform our understanding of language models and is connected to several other existing theoretical results. As an illustration of the usage of our framework, we derive a bound on in-context learning loss with our framework. Finally, we support our framework with empirical experiments and provide possible future research directions.
Pelican Soup Framework:
A Theoretical Framework for Language Model Capabilities
Ting-Rui Chiang University of Southern California tingruic@usc.edu          Dani Yogatama University of Southern California yogatama@usc.edu
1 Introduction
Large language models (LLMs) have demonstrated the capability to perform downstream natural language processing (NLP) tasks. By following instructions, LLMs can perform tasks with zero-shot examples, demonstrating their reasoning capability. With some input-output examples provided in the prompt, LLMs can also perform tasks without instructions, which is known as in-context learning (ICL) (Chowdhery et al., 2022). In particular, Brown et al. (2020) show that LLMs can perform ICL for classification tasks even when the verbalizers (labels present in the demonstration) are semantically irrelevant to the task, e.g., foo/bar instead of negative/positive Wei et al. (2023). However, it is unclear how pretraining leads to these capabilities.
To explain how LLMs acquire these capabilities, we propose a simple theoretical framework, the Pelican Soup framework in §2. Our framework is based on some very general assumptions, such as the logical consistency of training documents (Assumption 3.2) and the variable meanings with which a phrase can be associated in different contexts (Assumption 3.3). Our framework also includes a simple formalism for NLP tasks, which helps explain why LMs can follow instructions.
In §4, we show how we can use this framework to analyze LLM’s ICL capability, which mitigates the limitations of previous theoretical analyses. For example, in the first theoretical analyses of ICL, Xie et al. (2022) assumes that the general text for training LMs is from a hidden Markov model (HMM), which may be an oversimplification of natural language. In comparison, our framework does not require this strong assumption.
Our framework offers unique insights compared to existing theories. For example, although the generation process by Zhang et al. (2023) is more general than the HMM assumption by Xie et al. (2022), it lacks groundings in real-world linguistic phenomena. Our framework mitigates this limitation, as it helps us better explain the physical meaning of the terms in the bound on the ICL loss and shows how the terms reflect real-world practices, such as the choice of verbalizers and the distribution of test examples. Compared to the work by Hahn and Goyal (2023), our theory allows us to bound ICL loss without relying assumptions on the grammar that generate the data.
Furthermore, in §5, inspired by the cognitive science theories Fodor (1975, 2008); Piantadosi (2021), early development of artificial intelligence  (Siskind, 1996; Murphy, 2004) and formal linguistics Carnap and others (1968); Bresnan and Bresnan (1982); Steedman (1987, 1996); Sag et al. (1999), we provide an extension of our framework to explain why generalization is possible. The extension also connects our framework to other theoretical results. For example, our extension instantiates the complex skills in the theory by Arora and Goyal (2023). In §6, with an additional assumption, we can relate the ICL bound to the description length of the underlying input-output mapping function as done by Hahn and Goyal (2023).
Our work informs future LLM research directions. Scientifically, we shed light on how properties of the training data contribute to the ICL and the instruction-following capabilities of LLMs. The framework also shows how linguistic and psychology studies can inform our understanding of modern NLP. Practically, we highlight the importance of acquiring knowledge about the interrelation between concepts through pretraining. As shown in previous studies, the language modeling objective is inefficient for knowledge acquisition Allen-Zhu and Li (2023); Chiang et al. (2024). We suggest developing a better pretraining technique is crucial for future NLP development. Our proposed experimental setups can also facilitate future studies on LLMs’ acquisition of the capabilities.
2 The Pelican Soup Framework
We aim our theoretical framework at explaining why LLMs can perform well on prompts for downstream tasks even though the prompts have a different distribution than the training corpus. Therefore, our framework includes assumptions qualifying the training corpus distribution in (§3.1) and a formalism for NLP tasks (§3.2). Later, we will show how this framework allows us to bound the loss of ICL.
3 Motivation
The Pelican Soup game inspires our framework. It is a game involving a puzzle master who has a story in their mind. The game participants ask the puzzle master questions. The goal of the participants is to recover the story based on the answer to the yes/no questions. An observation is that once the participants recover the story, they can answer any questions about the story. This is similar to ICL, where the model is required to “figure out” the underlying task based on the input-output pairs in the given demonstration. Once the model identifies the underlying task, it will be able to predict the label of any input. Therefore, the story has a similar role as a latent variable defining the task, and the yes/no questions are similar to the demonstrations for ICL. We include an example in §A.
Given the above observation, we can study ICL by considering why humans can solve Pelican Soup riddles. We conjecture that this is because the person who makes the story and the ones who solve the riddle share the same (or similar) commonsense (McCarthy, 1960) about logical relationships among the events in this world (Schank and Abelson, 1988). Based on this shared commonsense, the game participants can eliminate hypothesis contradictory to the information that the asked yes/no questions reveal, which allows the participants to uncover the underlying story. This idea of “shared commonsense” inspires us to introduce the notion of “no ambiguity” and “consistency” in our framework.
3.1 Training Data Distribution
Our theory framework is based on the interrelations between the semantics of sentences. Thus, we first make a general assumption:
Assumption 3.1 (No ambiguity).
For all pairs of sentences $x_{1},x_{2}$ in a language, we assume humans can determine the relationship between $x_{1}$ and $x_{2}$ is contradiction, entailment or neutral.
Assumption 3.1 allows us to qualify sentences that can cooccur in a paragraph with non-zero probability. That is, a paragraph generally does not contain self-contradictory information:
Assumption 3.2 (Consistency).
Any paragraph with non-zero probability mass does not contain two sentences contradicts to each other.
To show how modeling natural language leads to the ICL capability, we further introduce the notion of expression-meaning association as a latent variable. It reflects the fact that language allows us to associate meanings with the surface form expressions quite freely. For example, when “she” or the human name “Emily” is present in a paragraph, it is associated with a certain person of certain characteristics, which exhibit its meaning.
Meanwhile, the usage of the expression depends on the meaning it is associated with and is consistent within its context. For example, if “she” is associated with the sentence “a person who has a house”, then by Assumption 3.2, the sentence “she has no property” will have 0 probability mass. Moreover, when we want to refer to “the person who has a house” instead of repeating the sentence again, we use “she” as an abbreviation.
For simplicity, we only consider single-word expressions and assume that such association is consistent throughout a document.
Assumption 3.3 (Expression-meaning association).
There is a set of words $\Gamma$ such that for every document in the training data, some $r\in\Gamma$ in the document is associated with a meaning represented as a set of sentences $Z_{r}$ with a prior distribution $\Pr(Z_{r})$. Any $z\in Z_{r}$ present in the document can be replaced with $r$ without breaking the logical consistency of the document.
Adjectives such as “good” and “bad” are also expressions that can be associated with variable meanings, and their meanings also depend on the context. However, the association may not be as variable as pronouns’. We reflect this with a prior distribution for the meaning with which an expression is associated later in our theoretical analysis.
Finally, we assume a document is a set of paragraphs where some expressions in $\Gamma$ are present:
Assumption 3.4 (Document).
A document is a concatenation of paragraphs containing $r\in\Gamma$ separated with a delimiter $d$ (e.g., a blank line).
Based on the above assumptions, we can define a perfect LM.
Definition 3.5 (Perfect LM).
We define an LM to be perfect if it only generates outputs satisfying the above assumptions.
In the following, we will show that a perfect LM can follow instruction and do ICL.
3.2 A Formalism for NLP Tasks
With Assumption 3.1, we propose a simple formalism: For any objective or prescriptive NLP task (Rottger et al., 2022) that maps an input $x$ to a set of acceptable outputs $Y$, that task can be described with some task instructions $u$ such that $u\wedge x\wedge y$ does not cause a contradiction if and only if $y\in Y$.
For example, if the task is a generation task, such as solving a math word problem, the instruction $u$ may specify the format (e.g., “Think step-by-step and output the answer after ####”) and the input $x$ is a math word problem to be solve. LLMs’ response is an acceptable answer if and only if the response does not cause a contradiction to $u$ and $x$, i.e., it follows the instruction and the reasoning process aligns with the problem $x$. When it is a classification task, the instruction $u$ should contain descriptions about each label $y$. For example, we can formulate the sentiment analysis task over movie reviews as $u=\langle v_{+},v_{-}\rangle=\langle\text{``I like the movie''},\text{``I dislike the movie''}\rangle$.
Under this formalism, it is trivial that perfect LLMs can follow instructions and solve tasks. It is because when the prompt for an LLM is the concatenation of the instruction and the task input, $u;x$, Assumption 3.2 (the consistency assumption) ensures that a perfect LLM only generates $y$ such that $y$ is logically consistent to its prompt $u;x$. Based on our NLP task formulation, if $y$ is logically consistent to $u$ and $x$, then $y$ is in $Y$, meaning that $y$ is an acceptable answer.
Two intricate questions remain, how can LLMs perform ICL? and why is it possible for an LM to generalize to unseen instructions? We discuss these two questions in §4 and §5.
4 Bounding ICL Loss
To see how we can analyze ICL with our framework, we first setup a latent variable model with our framework using Assumption 3.3 (the expression-meaning association assumption). In this latent variable model, the latent variables are the association between the expressions in the document and their meanings. Assumption 3.2 (the consistency assumption) defines the support of the distribution of the latent variable conditioning on a context: Associations that break the logical consistency of the context have zero probability. Assumption 3.2 also implies that, give a prefix, continuations that imply invalid associations have zero probability mass.
We analyzing ICL with this latent variable model by focusing on the latent variable for the association between the verbalizers and the meanings. Given a classification task, the underlying latent variable value $z^{*}$ that produces the demonstration is the association where verbalizers are associated with the class descriptions of the task (following the definition of tasks in §3.2). For example, when “foo” is used to represent the positive class of an sentiment classification task, “foo” should be associated with the description of the positive class, i.e., “I like the movie”. If an LLM can infer this underlying association $z^{*}$ based on the examples in the demostration and output a continuation satisfying Assumption 3.2, then it can predict the correct verbalizer for a test example.
Formally, adapting and combining the analyses by Zhang et al. (2023) and Hahn and Goyal (2023), we have the following theorem.
Theorem 4.1 (Average ICL Likelihood).
Denote a sequence of input-output pairs as $S_{t}=x_{1},r_{1},d,x_{2},r_{2},\cdots,x_{t},r_{t},d$, where $r_{i}$ is the correct verbalizer with which the label of $x_{i}$ is associated for $i=1,2,\cdots,t$ and $d$ is the delimiter that separates the examples. Let the description of a task that maps inputs to classes $\mathcal{Y}$ be $\{v_{y}\}_{y\in\mathcal{Y}}$ and $z^{*}$ represents that the task descriptions $\{v_{y}\}_{y\in\mathcal{Y}}$ are associated with the corresponding verbalizers $\{r_{y}\}_{y\in\mathcal{Y}}\subset\Gamma$ used for ICL. We have for any $T\in\mathbf{Z}^{+}$, the average log likelihood that the LM predicts the correct verbalizer $r_{t}$ for $t=1,2,\cdots,T$ is bounded as:
$\begin{split}&\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1})\\ &\geq\begin{aligned} &\frac{1}{T}\log\Pr(z^{*})\\ &+\frac{1}{T}\sum_{t=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},S_{t-1})\\ &+\frac{1}{T}\sum_{t=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1})}{\Pr(x_{t}|S_{t-1})}\end{aligned}\end{split}$ (1)
When the last two terms on the right-hand side are nonnegative, Eq. 1 shows that the average log likelihood converges to 0 in $\mathcal{O}(1/T)$. We discuss the terms on the right-hand side below.
The second term is 0. This is because Assumption 3.2 ensures that the probability $\Pr(r|x_{t})$ is zero if $r$ is not an acceptable output for $x_{t}$ (i.e., when $r\neq r_{t}$ for classification tasks).
We then look at the last term. This term is 0 when $x_{t}$ is conditionally independent of $z^{*}$ as assumed by Zhang et al. (2023). However, this may be an oversimplification because, in natural language, the transition from $x_{t}$ to its next token depends on the content of $x_{t}$. Fortunately, this assumption may actually be unnecessary for convergence because $x_{t}$ is an example from a downstream task related to $z^{*}$; it is likely that
$\Pr(x_{t}|z^{*},S_{t-1})\geq\Pr(x_{t}|S_{t-1}),$
which implies that this term is non-negative, and we can thus ignore this term. More rigorously, we can show the following corollary.
Corollary 4.2 (Expected Average ICL Loss).
Assuming $z^{*}$ is the association suggesting that the sub-sequences separated by the delimiter $d$ are independent of each other, namely, $\Pr(x_{1},r_{1},d,x_{2},r_{2},d,\cdots,x_{t},r_{t},d|z^{*})=\prod_{t=1}^{T}\Pr(x_{t},r_{t},d|z^{*})$. If the downstream task data distribution $\mathcal{D}_{X}$ follow $\Pr(x|z^{*})$, then we can bound the average ICL likelihood over the distribution as:
$\begin{split}&\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\left[\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1})\right]\\ &\geq\frac{1}{T}\log\Pr(z^{*}).\end{split}$ (2)
The right-hand side of Eq. 2 characterizes the convergence rate and reflects the difficulty of doing ICL. It reflects that when the association between label description and verbalizer is uncommon in the training data (e.g., associating “positive” with “This movie is bad.”), doing ICL is more difficult.
Note that we can extend the results to generation tasks. For generation tasks, we usually use a separator (or a short text span) between $x_{t}$ and $r_{t}$. We can see the separator as an expression that can be associated with different meanings, so the latent space for $z$ is the meaning the separator can be associated with and $z^{*}$ is the association between the separator and the task instruction (details in §C).
5 Generalization
In §2, we assume a latent model, which, however, poses a dilemma. Language can encode various meanings. If the sequence length is unconstrained, the corresponding semantic space can even be infinite. However, if the latent space is infinite, the limited training data would not cover the entire latent space. Without any assumption on the latent space (e.g., the relation between the states in the space), it is impossible to discuss the generalization to unseen latent states. Thus, we provide an extension of our theoretical framework.
Assumption 5.1 (Meaning representation).
There exists (1) a finite set of atom concepts $\Sigma$, (2) a knowledge base $\mathrm{KB}$ consisting of logical rules between the atom concepts in $\Sigma$, and (3) a function $f$ that can map any sentence in language to its meaning represented as a logical formula with operands in $\Sigma$ such that for any two sentences $s_{1},s_{2}$, the logical relation between $s_{1}$ and $s_{2}$ judged by humans is the same as $f(s_{1})$ and $f(s_{2})$ given the rules in the knowledge base $\mathrm{KB}$.
The three elements of this assumption correspond to theories in various fields. The notion of atom concepts aligns with cognitive psychology studies that hypothesize the existence of a set of mental tokens Fodor (1975, 2008) and a recent study Piantadosi (2021) suggesting semantics can be encoded with the combination of only a few symbols. The notion of knowledge base follows the early formulation of AI (Siskind, 1996; Murphy, 2004). As for the existence of a parsing function $f$, it follows the long history of linguistics studying the relationships between natural and formal languages (Carnap and others, 1968; Bresnan and Bresnan, 1982; Steedman, 1987, 1996; Sag et al., 1999; Frege and others, 1879; Peirce, 1883).
This assumption suggests that if we have the parsing function $f$, solving NLP tasks only requires a finite-length program that can do logical reasoning by manipulating logical symbols according to logical induction rules. If a model can learn this program, then it can perform a task even if this task is not in the training data. This assumption of a finite $\Sigma$ also instantiates the concept of “language skills” by Arora and Goyal (2023), and their theoretical results are thus applicable.
Figure 1: Calcutec examples for training, in-context learning (ICL), and chain-of-thought (CoT).
6 Relating to Description Length
We can see $\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$ as the difficulty of the example by having an additional assumption:
Assumption 6.1.
In some documents in the training data, the paragraphs are constituted with steps in a logical induction process, with some steps randomly dropped.
This kind of document may be prevalent in training data. Essays arguing some claims are one example. To be convincing, these essays should proceed like a proving process that induces their conclusions. Documents describing a series of events can be another example, as events follow commonsense and develop progressively.
With this assumption and some regularity assumptions on the data distribution, we can have
$\Pr(r_{t},d|x_{t},z^{*},S_{t-1})\leq c\cdot\ell(x_{t}),$ (3)
where $\ell(x_{t})$ is the number of reasoning steps required to solve the task, and $c$ is a constant. This $\ell(x_{t})$ corresponds to the description length of the function that maps the inputs to their label in the loss bound by Hahn and Goyal (2023) (more discussion in Appendix E).
7 Inspecting Generalization Empirically
Although we have shown that perfect LMs can follow instructions (§3.2) and perform ICL (§4) when training data satisfy the assumptions in §2, it remains unverified whether imperfect real-world LMs exhibit the same capabilities. For LMs to demonstrate these abilities, they must generalize well from the training set to the prompts used during inference. However, analyzing how deep models generalize inherently requires additional assumptions about the model, which is beyond the scope of this paper. Thus, to address this gap, we leverage our extended framework (§5) to characterize distribution shifts and design experiments that empirically examine the generalization of LM.
7.1 Inspecting the ICL Capability
Firstly, we present a synthetic setup, Calcutec, as a concrete instantiation of our theoretical framework. With Calcutec, we show that Transformers can acquire ICL capability by modeling the linguistic characteristics specified in our framework.
7.1.1 Calcutec
Setup
Following the extended framework in §5, we construct a pseudo-language:
  • Logic model: We use a subset of propositional logic as our logic model. We only consider Horn clauses (Horn, 1951), i.e., formulas in the form $A\wedge B\to C$.
  • Atom concepts: We use 100 symbols as our set of atom concepts $\Sigma$.
[…]
  • We have a set $\Gamma=\{r_{i}\}_{i=1}^{4}$ representing the expressions described in Assumption 3.3.
Training Dataset.
Following Assumption 3.4, a document is a concatenation of paragraphs separated by delimiter “;” and ends with “.”. In our synthetic language model training dataset, each document contains 16 paragraphs.
Following Assumption 3.2, each paragraph consists of sentences logically coherent to each other. Because sentences in the real world are not ordered arbitrarily, we follow Assumption 6.1 and generate random paragraphs following the structure of logical proofs. Each paragraph represents an induction process of $P\models g$ for some randomly selected $P\subset\Sigma$ and $g\in\Sigma$. Each sentence in the paragraph is a sentence representing a reasoning step for $P\models g$. We separate the sentences in the sequence by commas. To simulate the noise in the real world, we further apply perturbations that skip some steps with a skip rate $p_{skip}$.  [^1] Models can acquire in-context learning ability even with $p_{skip}=0$ (Figure 6 in the appendix).
Following Assumption 3.3, after we generate a document, we randomly associate some symbols $A,B\subset\Sigma$ with $r_{a},r_{b}\in\Gamma$ respectively. We reflect this association in the generated document by replacing symbols in AA and $B$ with expression $r_{a},r_{b}\in\Gamma$ respectively (details in Appendix F.1).
Downstream Tasks.
Following the formalism in §3.2, we define a binary classification task by defining the descriptions $v_{+}$ and $v_{-}$ of the positive and negative classes, respectively. We use the disjunctions of atom concepts (i.e. in the form of $a_{1}\vee a_{2}\vee\cdots$) as descriptions of classes. We create five downstream tasks using different disjunctions. Each input is a subset of variables in $\Sigma$ from which we ensure only one of the classes can be induced.
Demonstration.
We represent an input-label pair as $x_{1}x_{2}\cdots\to r$, where $x_{1}x_{2}\cdots$ is the input part and $r\in\{r_{+},r_{-}\}\subset\Gamma$ is an expression in $\Gamma$ serving as the verbalizer.
Chain-of-thought.
A chain-of-thought is in the same format as the training data but ends with an expression $r\in\{r_{+},r_{-}\}$, e.g., $x_{1}x_{2}\cdots\to x_{3};x_{3}\cdots x_{4}\to r_{+}$. This chain-of-thought reflects the step-by-step induction process from the inputs to the label. We show an example in Figure 1.
7.1.2 Distribution Shifts
We make experimental designs to simulate the real-world distribution shifts from training to inference:
Format Mismatch.
The reasoning steps are present in the training data but not in the prompts.
Verbalizer Mismatch.
When we choose the expressions in $\Gamma$, we assign the probability mass 45%, 45%, 5%, 5% to $r_{1},r_{2},r_{3},r_{4}$. In this way, we can inspect the effect of using less frequent verbalizers.
Unseen Tasks.
To investigate whether the model can generalize to a new combination of formulas unseen in the training data when we generate our training data, we ensure that the expressions in $\Gamma$ are only associated with the disjunctions of two atom concepts $s_{1},s_{2}$ from a strict subset of all possible combinations $\Sigma\times\Sigma$. We then test the trained model on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of the unseen combinations. We also test the models on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of three atom concepts $\in\Sigma\times\Sigma\times\Sigma$.
7.1.3 Experiment Details
We use $p_{skip}=0.25$ in our experiment, generating 60,000 documents with 16 paragraphs (as described above). Among them, we use 10k for validation. We train a 6-layer Transformer (Vaswani et al., 2017) model until the loss on the validation set converges using the standard autoregressive loss. We include additional setups in §F.3.
7.1.4 Results and Discussion
$r_{1},r_{2}$ $r_{3},r_{4}$
Task ICL CoT ICL CoT
Single 57.1 91.7 55.6 92.0
Double 53.5 76.3 51.1 77.1
Triple 53.0 73.0 51.7 73.4
Table 1: The 4-shot accuracy of ICL versus chain-of-thought (CoT) using different verbalizers.
Figure 2 shows that LMs trained with Calcutec can perform in-context learning. This evidence supports our Pelican Soup framework and aligns with our theoretical analysis in §4: The ICL accuracy follows a trend $\exp(c/T)$ for some $c<0$, as our theoretical result in Eq. 1.
We further inspect the ICL performance under the distribution shifts described in §7.1.2. For the infrequent verbalizer, we observe similar performance regardless of the frequency of the verbalizers ($r_{1},r_{2}$ versus $r_{3},r_{4}$). For the unseen tasks, Figure 2 shows that the model has similar performance in tasks defined with seen and unseen combinations of atom concepts (dotted and solid lines), even generalizing to tasks defined with three latent concepts (green lines). In sum, the results show that the model can generalize well under several distributional shifts.
(a) Verbalizers = $r_{1},r_{2}$
(b) Verbalizers = $r_{3},r_{4}$
Figure 2: In-context learning accuracy with Calcutec when using different verbalizers ($r_{1},r_{2}$ or $r_{3},r_{4}$). The dotted lines represent the performance of unseen combinations described in §7.1.2. The colors represent the number of atom concepts each class ($v_{+}$ or $v_{-}$) is associated with. The main lines represent the average accuracy of 5 tasks. Lines in the lighter color represent the individual tasks.
We also experiment with 4-shot learning using chain-of-thought. Table 1 shows that the model also benefits from chain-of-thought. We conjecture that it is because chain-of-thought has a format more similar to the format for training.
7.2 Inspect Instruction-following Capability
In addition to ICL, we inspect another intriguing capability of LLMs, the capability of following unseen instructions. Our extended framework in §5 suggests that we can see following unseen instructions as doing a sentence completion task that involves atom concepts (or say “skills” in Arora and Goyal (2023)) whose composition is different from the ones involved when completing sentences in training data. We also suggest that LMs acquire the capability of following these unseen instructions by modeling the interrelations between the atom concepts involved. To show the plausibility of this, we experiment with an arithmetic task.
7.2.1 Arithmetic Task
Single-step 0 + 1 * 2 + 3 * 4 = 10 .
Multi-step 0 + 1 * 2 = 2 . 3 * 4 = 12 . 2 + 12 = 14 .
Test - seen 2 * 3 + 4 * 5 + 6 =
Test - unseen 2 + 10 + 4 * 11 * 6 =
Table 2: Training (first two rows) and test (last two rows) examples used in the arithmetic task §7.2.1.
We utilize the algebraic structure of the integers under modulo addition/multiplication to construct a language where each expression is constantly associated with a meaning and the atom concepts include 0 to 15, the symbol “+”, “*”, “=”, “.”. Each sentence in this language is in the form of an equation with additions and multiplications.
We generate LM training data in a multi-step and a single-step setup. In the multi-step setup, each example contains the steps required to compute the result, each of which involves two to five numbers (details in §G). We expect that LMs can learn the interrelation between atom concepts by modeling these steps. In the single-step setup, the final result is computed in a single step.
To simulate the scenario where completing the prompts in the test set involves a different distribution of atom concepts from that of the training examples, we ensure that the five numbers for each training example are either all from $\{0,\cdots,9\}$ or all from $\{6,\cdots 15\}$. We then evaluate the model with a seen and an unseen setups. In the seen setup, we prompt the model with 5 numbers sampled in the same way as the training set. In the unseen setup, we prompt the model to complete equations with 5 numbers where the first, third, and fifth numbers are from $\{0,\cdots,5\}$ and the other two numbers are from $\{10,\cdots,15\}$ (examples in Table 2).
7.2.2 Results and Discussion
test $\backslash$ train Multi-step Single-step
Seen 97.5 (0.3) 99.3 (0.6)
Unseen 75.3 (2.0) 49.7 (1.0)
Table 3: The average accuracy of LMs trained for the arithmetic tasks with 5 random seeds.
Table 3 shows that all LMs achieve a near-perfect accuracy in the seen setup but the LMs trained with the multi-step training set achieve significantly better accuracy in the unseen setup. This indicates that modeling the multiple steps in the training data may allow the model to learn the interrelation between the atom concepts and generalize to prompts that involve unseen compositions of atom concepts to some extent, as suggested by our extended framework in §5. This is also aligned with the success of symbolic chain-of-thought distillation (Li et al., 2023a; Hsieh et al., 2023; Shridhar et al., 2023).
task SST-2 CR MR Subj
direct 63.0 61.7 59.2 51.0
direct w/ foo/bar 51.7 53.7 52.2 57.1
pronoun 65.3 62.9 56.7 62.2
Table 4: The accuracy of using task-specific templates/verbalizers (direct) (Min et al., 2022a) v.s. using task-agnostic templates/pronouns for 16-shot in-context learning with GPT2-Large.
8 Real-world Evidence
We inspect how likely that LMs learn ICL by modeling the expression-meaning association as our framework suggests. We specifically inspect a small model GPT2-Large and the setting in which we use pronouns as verbalizers. Because pronouns are reference words frequently associated with different meanings, we expect that even a small model can do ICL well with pronouns. To do so, we experiment with the template [input]”, [verbalizer] thought. and use “he”, “she” as verbalizers. We follow the setup in Min et al. (2022a) and compare the accuracy of binary classification tasks, including SST-2 Socher et al. (2013), CR Hu and Liu (2004), MR Pang and Lee (2005), and Subj Pang and Lee (2004), using GPT2-Large.
Table 4 shows that this task-agnostic template with pronouns is competitive with those task-specific templates. This aligns with our conjecture that the high frequency of pronouns in the training set allows smaller LMs to acquire the ICL capability. It also shows that, unlike what Wei et al. (2023) claims, not only larger models can do in-context learning with task-irrelevant verbalizers. On the other hand, using task-irrelevant verbalizers “foo” and “bar” has lower performance (which aligns with the observation of Wei et al. (2023)). This may be because learning to perform ICL with those low-frequency words requires more training data, and thus a larger model size.
9 Related Work
Since Brown et al. (2020) discovered large language models’ in-context learning ability, some theoretical works have attempted to explain how language models acquire this ability. Based on a hidden Markov model (HMM) assumption on the language generation process, Xie et al. (2022) suggested that in-context learning is an implicit Bayesian inference process. Hahn and Goyal (2023) defined the generation process with Compositional Attribute Grammar, which is weaker than the HMM assumption, explaining the in-context learning ability with the minimum description length. They also studied the compositionality of natural language tasks with function compositions. Zhang et al. (2023) assumed a more general latent variable model. Arora and Goyal (2023) analyze the emergence of skills based on the scaling law (Hoffmann et al., 2022). While their analysis assumes a set of atomic skills for NLP tasks, our framework is based on a set of atom concepts.
There were also many empirical studies on the in-context learning ability. Some works focused on the effect of the instruction (Webson and Pavlick, 2022; Lampinen et al., 2022; Jang et al., 2023), while some focused on the examples in the demonstration (Liu et al., 2022; Lu et al., 2022; Sorensen et al., 2022; Min et al., 2022b; Yoo et al., 2022; Ye et al., 2023; Chang and Jia, 2023; Ye et al., 2023; Wang et al., 2023b; Kossen et al., 2023). Shin et al. (2022) found that not all training corpora led to in-context learning ability. Prystawski and Goodman (2023) used synthetic data to suggest that the pretraining dataset’s locality structure contributes to the reasoning steps’ effectiveness. Wang et al. (2023a) studied the reasoning steps in chain-of-thought. Akyürek et al. (2024) formulated ICL as learning a formal language from demonstrations and benchmarked model families.
Some previous work studied in-context learning as a meta-learning-like problem (Chen et al., 2022). Some works focused on the relationships between in-context learning and optimization algorithms (Garg et al., 2022; von Oswald et al., 2022; Akyürek et al., 2023; Fu et al., 2023; Guo et al., 2023). Some works inspected the mechanism of ICL in transformer models (Hendel et al., 2023; Bietti et al., 2023; Todd et al., 2023; Shen et al., 2023; Bai et al., 2023). Chan et al. (2022) studied the properties of dataset distribution that could contribute to the in-context learning ability. Li et al. (2023b) provided generalization bounds based on the stability of Transformer models and the distance of downstream tasks. We instead focus on how the pretraining data in natural language contributes to the ICL learning ability.
10 Conclusion and Future Work
In this work, we propose a framework that explains how linguistic phenomena in the training corpus lead to LLMs’ ICL and instruction-following capability. Compared with previous works Xie et al. (2022); Zhang et al. (2023), our latent model better reflects the complexity of language. By introducing the notion of knowledge base and logic system, our framework provides insights into how LLMs can generalize from pretraining to downstream tasks, instantiating a setup compatible with the assumptions made by Arora and Goyal (2023). We also relate our bound to the function description length discussed by Hahn and Goyal (2023).
Our framework illuminates a few possible directions for improving LLMs:
  1. 1.
    Our work highlights the importance of learning the interrelation between meanings. As previous works have shown that the language modeling objective is inefficient for this purpose (Allen-Zhu and Li, 2023; Chiang et al., 2024), we suggest that developing a more sophisticated learning algorithm is crucial.
  2. 2.
    Our theory illustrates how linguistic studies on parsing functions can inform LLM research, offering a new theoretical foundation for analyzing LLM generalization.
  3. 3.
    The experimental results of our arithmetic task show that Transformer models can generalize to unseen prompts by modeling the intermediate step-by-step reasoning process. This may be related to the success of the symbolic chain-of-thought distillation (Li et al., 2023a; Hsieh et al., 2023; Shridhar et al., 2023). Investigating and strengthening the mechanism can improve the efficiency of LM training.
11 Limitations
A limitation of our framework is that, as most theoretical studies do, we simplify the real-world scenario to draw insights. One simplification we make is that, we do not take the noise in LLMs’ training data into account. As we may need to make more assumption on the noise to establish a generic theoretical result, we leave it for future study. Another simplification is that, we assume that the language model can perfectly model the distribution of natural language (i.e., only generates outputs satisfying the assumptions in §3.1, especially Assumption 3.2). However, it is unlikely to be the case in practice. On the one hand, the training data may not cover all the test cases. On the other hand, LLMs may not perfectly generalize from the training set. We need to make more assumptions on the training/test data distribution and/or have a deeper understanding on how deep learning models generalize to alleviate this assumption. Therefore, we deem this out of the scope of this paper.
We also note some limitations of our experiments. As commonly seen in theoretical work, our experiment setup in §7.1 is a simplification of natural language. Our intention is to isolate and illustrate specific theoretical mechanisms without confounding variables introduced by the full complexity of natural language. The experiment in §8 shows the correlation between better ICL performance and the use of pronouns as verbalizer. Although this serves as evidence aligning with our theoretical framework, more studies are required to establish the causal relationship.
Appendix A A Pelican Soup Example
We include an example of a Pelican Soup game:
Puzzle master: A men walks into a restaurant and orders pelican soup. After taking a sip, he loses his mind. Why?
Participants: Is it because the soup is not cooked well?
Puzzle master: No.
Participants: Is it because the soup toxic?
Puzzle master: No.
Participants: Does the soup remind him something?
Puzzle master: Yes.
Participants: Did someone cook pelican soup for him?
Puzzle master: Yes.
Participants: Is that person still alive?
Puzzle master: No.
For the sake of aesthetics, we do not include the latent story here. If you are interested, please check it online.
Appendix B Proof of Theorem 1
Let $S_{t}=x_{1},r_{2},d,x_{2},r_{2},d\cdots,x_{t},r_{t},d$.
$\Pr(z|S_{t})$
$=$ $\frac{\Pr(S_{t}|z)\Pr(z)}{\sum_{z}\Pr(S_{t}|z)\Pr(z)}$
$=$ $\frac{\Pr(z)\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z,S_{i-1})}{\sum_{z^{\prime}}\Pr(z^{\prime})\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z^{\prime},S_{i-1})}$
$P(x_{t+1},r_{t+1},d|S_{t})$
$=$ $\sum_{z}\Pr(x_{t+1},r_{t+1},d|z,S_{t})\Pr(z|S_{t})$
$=$ $\frac{\sum_{z}\Pr(z)\prod_{i=1}^{t+1}\Pr(x_{i},r_{i},d|z,S_{i-1})}{\sum_{z^{\prime}}\Pr(z)\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z^{\prime},S_{i-1})}.$
Thus, it holds that
$-\sum_{t=0}^{T}\log\Pr(x_{t+1},r_{t+1},d|S_{t})$
$=$ $\begin{aligned} &-\sum_{t=0}^{T}\Big(\\ &\log\sum_{z}\Pr(z)\prod_{i=1}^{t+1}\Pr(x_{i},r_{i},d|z,S_{i-1})\\ &-\log\sum_{z}\Pr(z)\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z,S_{i-1})\\ &\Big)\end{aligned}$
$=$ $-\log\sum_{z}\Pr(z|K)\prod_{i=1}^{T+1}\Pr(x_{i},r_{t},d|z,S_{i-1})$
$\leq$ $-\log\Pr(z^{*})\prod_{i=1}^{T+1}\Pr(x_{i},r_{t},d|z^{*},S_{i-1})$
$=$ $-\log\Pr(z^{*})$
$-\sum_{i=1}^{T+1}\log\Pr(x_{i},r_{i},d|z^{*},S_{i-1})$
$=$ $-\log\Pr(z^{*})$
$-\sum_{i=1}^{T}\log\Pr(r_{i},d|x_{i},z^{*},S_{i-1})$
$-\sum_{i=1}^{T}\log\Pr(x_{i}|z^{*},S_{i-1}).$
Thus,
$-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1})$
$\leq$ $-\frac{1}{T}\bigg(\log\Pr(z^{*})$
$+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$
$+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1})}{\Pr(x_{t}|S_{t-1})})$
$+\frac{1}{T}\sum_{i=1}^{T}\log\Pr(d|r_{t},x_{t},S_{t-1}\bigg).$
$\leq$ $-\frac{1}{T}\bigg(\log\Pr(z^{*})$
$+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$
$+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1})}{\Pr(x_{t}|S_{t-1})}\bigg)$
Appendix C Bounding ICL for Generation
We can extend the ICL bound shown in Theorem 1 to generation tasks. When using ICL for generation, in the demonstration part of the prompt, people usually include a separator between the input and the output. For example, if the task is to translate a sentence to English, people may write the demonstration in the format “[input] should be converted to [output].” This “should be converted to” is an expression that can be associated to the “meaning” of the task, namely “should be mapped to a translation in English.” The intuition of the theorem is that, a language model may be able to uncover this association through the demonstration examples via the same mechanism as how it perform ICL for classification tasks.
Theorem C.1 (Average ICL Loss for Generation).
Let $z^{*}$ represent the association between a separator $\xi$ and a task descriptions such that for any input $x$, $P(x,\xi,r|z^{*})>0$ only if $r$ is one of the correct outputs as specified by the task. Let $K$ be the constraints used for decoding, and $\dot{g}$ be the event where a document follows certain formats. Let $R_{t}$ be the set of correct outputs for $x_{t}$, $S_{t}=\{x_{1},\xi,r_{2},d,x_{2},\xi,r_{2},\cdots,x_{t},\xi,r_{t},d|r_{1}\in R_{1},r_{2}\in R_{2},\cdots,r_{t}\in R_{t}\}$. We have for any integer $T>0$, the average cross-entropy loss of ICL is bounded as:
$\begin{split}&-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(R_{t}|x_{t},\xi,S_{t-1})\\ &\leq\begin{aligned} &-\frac{1}{T}\log\Pr(z^{*})\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\Pr(R_{t},d|x_{t},\xi,z^{*},S_{t-1})\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\frac{\Pr(x_{t},\xi|z^{*},S_{t-1})}{\Pr(x_{t},\xi|S_{t-1})}\end{aligned}\end{split}$ (4)
Proof.
If we see $\xi$ as a part of the $x$’s in the proof for Theorem 1, then following the steps in §B, we can have Eq. 4. ∎
Appendix D Proof of Corollary 2
The second term in the right-hand side of Eq. 1 is zero when the decoding constrain $K$ is imposed. Therefore, it suffices to prove the last term is non-negative in expectation.
$\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\sum_{x_{1},x_{2},\cdots,x_{T}}\Pr(x_{t}|z^{*},K)\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\mathrm{KLD}(\Pr(x_{t}|z^{*},K)||\Pr(x_{t}|S_{t-1},K))\geq 0$
Appendix E The Connection between $P(r_{t}|x_{t},z^{*})$ and Function Description Length by Hahn and Goyal (2023)
Firstly, we make some regularity assumptions: Given a step-by-step reasoning process $\pi=s_{1},s_{2},\cdots,s_{n}$ for the induction process of $P\models Q$, in the training data,
  1. 1.
    each step may be dropped independently to each other with probability $p_{drop}$.
  2. 2.
    $\Pr(s_{i}|P,s_{1},s_{2},\cdots,s_{i-1})>p_{min}$ for all $i\in[n]$.
We first show how we derive Eq. 3: Based on Assumption 6.1,
$\begin{split}&\Pr(r_{t}|x_{t},z^{*})\\ &=\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*})\Pr(\text{$\pi$ is dropped}),\end{split}$
where $\Pi$ is a set of token sequences representing reasoning steps that induce $r_{t}$ from $x_{t}$. Let $\pi^{*}$ be the shortest proof in $\Pi$, we have
$\log\Pr(r_{t}|x_{t},z^{*})$
$=$ $\log\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*})\Pr(\text{$\pi$ is dropped})$
$\geq$ $\log\Pr(\pi^{*},r_{t}|x_{t},z^{*})\Pr(\text{$\pi^{*}$ is dropped})$
$\geq$ $p_{min}\log\ell(\pi^{*})+p_{drop}\log\ell(\pi^{*}).$
Then we can discuss the connection between $\Pr(r_{t}|x_{t},z^{*},\ddot{g})$ and the function description length by Hahn and Goyal (2023). We can see the dropped reasoning steps in $\pi^{*}$ as the hidden (tree) structure that maps $x_{t}$ to $r_{t}$ as the derivation tree $\tau_{\phi}$ in the bound of Hahn and Goyal (2023). The length of the reasoning steps thus corresponds to the description length of the derivation tree $\mathrm{D}(\tau_{\phi})$.
A major difference between the bound by Hahn and Goyal (2023) and our bound is that their bound has $\mathrm{D}(\tau_{\phi})$ constant to $T$ while our bound has $\sum_{t}\log\Pr(r_{t}|x_{t},z^{*},\ddot{g})$, which potentially grows proportionally to $T$. The cause of this difference is that, Hahn and Goyal (2023) assumes a structure that repetitively applies a function mapping in a document, and the number of repetition is independent to the complexity of the function mapping. In comparison, our framework does not make this assumption.
Appendix F Details of the Calcutec Experiment
F.1 Generation Process of the LM Training Data in Calcutec
We generate a paragraph based on Assumption 3.3 in the following step:
  1. 1.
    We pick a symbol $s$ from the symbols associated with $r_{a}$ uniformly at random.
  2. 2.
    We randomly generate a proof for $\mathrm{KB},P\models g$, where $P\subset\Sigma$ is the premise and $g\in\Sigma$ is the goal of the proof. We ensure that this proof contains the topic $s$.
  3. 3.
    We convert the proof tree to a sequence of proving steps by traversing the proving tree in a topological order with ties broken randomly. Each node in the proof tree corresponds to a rule in $\mathrm{KB}$, so the resulting sequence of proving steps consists of horn clauses in the form $a_{1}a_{2}\to b$. We separate the clauses in the sequence with commas.
  4. 4.
    We rewrite the first step of the proving process to contain the premises of the proof. Specifically, we replace the antecedent in the first formula with the premise $P$. We find that this step is necessary to prevent the language model trained on it from hallucinating irrelevant variables randomly. It is important for our experiment for chain-of-thought, but is not necessary for language models to learn the in-context learning ability.
Sample $r_{a},r_{b}$ from $\{r_{1},r_{2},r_{3},r_{4}\}$ with probability 0.45, 0.45, 0.05, 0.05.
Sample topic $S=\{s_{1},s_{2}\}\subset\Sigma$.
Initialize a document $D$ with empty string.
for $p=1,2,\dots,n_{par}$ do
  while True do
   Sample $s\in S$.
   Sample a set $X\subset\Sigma$ such that $\bigwedge_{x\in X}x\models s$.
   Run the resolution algorithm to get the set $M=\{m|X\models m\}$.
   Find an extra premise $x^{\prime}$ that can increase the depth of deepest proof tree for $X\models m$.
   Run the resolution algorithm to get the set $M^{\prime}=\{m|X\cup\{x^{\prime}\}\models m\}$.
   if $|M^{\prime}|>\frac{|\Sigma|}{2}$ then
     Reject the sampled $X\cup\{x^{\prime}\}$. $\triangleright$ We don’t want a premise that entails everything.
     Restart the while loop.
   end if
   Sample a $g\in M^{\prime}$ such that the proof tree for $X^{\prime}\models g$ contains $s$ and its depth $d_{min}$. $\triangleright$ We use $d_{min}=4$ in our experiments.
   Do topological sort to flatten the proof tree and convert it into a string.
   Append the string to $D$.
  end while
end for
for $s\in S$ do
  $\leftarrow DD$.replace($s$, $r_{a}$)
end for
Let $S^{\prime}=\{s^{\prime}_{1},s^{\prime}_{2}\}\in\Sigma$ be the top-2 frequent non $r_{a}$ symbols in $D$.
for $s^{\prime}\in S^{\prime}$ do
  $\leftarrow DD$.replace($s^{\prime}$, $r_{b}$)
end for
Algorithm 1 Pseudo code for the generation process of a Calcutec document used for training.
F.2 Perturbations in Calcutec
We apply two types of perturbations over the reasoning steps in Calcutec described in §7.1:
  1. 1.
    Random merge: At probability $p_{merge}$, for every two consecutive clauses where the consequence of the first one is in the antecedents of the second one, say $a_{1}a_{2}\to b_{1}$ and $b_{1}a_{3}\to b_{2}$, we merge them into a single clause $a_{1}a_{2}a_{3}\to b_{2}$.
  2. 2.
    Random drop: Given a clause $a_{1}a_{2}\cdots a_{n}\to b$. We drop each of the antecedents $a\in\{a_{1},a_{2},\cdots a_{n}\}$ at probability $p_{drop}$. We apply this dropping to every clause in the proof except the first one to ensure that we do not drop the premises.
We use $p_{merge}=p_{drop}=p_{skip}$.
Additionally, when flattening the proof trees with topological sort, we break the ties randomly. We also randomize the order of the symbols in the antecedents.
(a) document (tokens)
(b) paragraph (tokens)
(c) paragraph (steps)
(d) first $z$ (steps)
(e) document (tokens)
(f) paragraph (tokens)
(g) paragraph (steps)
(h) first $z$ (steps)
Figure 3: The distribution of lengths and the first step in each paragraph where $z$ is the consequence in the Calcutec dataset. The first/second row are the statistics before/after some steps are randomly dropped.
F.3 Extra Setups
Unseen Inference Process.
Based on Assumption 6.1 and the formalism of NLP tasks in §3.2, input-label pairs of a downstream task corresponds to prefix-reference pairs in a paragraph. To examine whether the trained model can generalize well when the induction process for the label is different from the induction process for the pronoun in the training data, we generate a training dataset where all the pronouns are induced from the premise with a left-branching proof tree with a depth equal to 2 (Figure 4(a)), while the test data contains samples whose labels are induced from the input with balanced trees (Figure 4(b)).
Different Input Lengths.
For each downstream tasks, we experiment with examples with different lengths. When the inference process is branching, having input length equal to $4$ makes the proving tree deeper.
No perturbations.
As described in §F.2, we apply some random perturbations on the proving process. We also experiment with the setup where we do not apply any perturbations.
With/Without Rewriting the First Step.
As described in §F.1, we rewrite the first step of the proof. We also experiment with the setup where we do not rewrite the first step.
Model Size.
We also experiment with different models sizes. We experiment with GPT-2 models that have 3, 4 and 5 layers.
F.3.1 Results and Discussion
Unseen Inference Process.
Figure 5(a) and Figure 5(d) show that the ICL performance on the branching examples is similar to the performance on the branching examples. It suggests that the model can generalize to examples that requires an unseen reasoning process. Interestingly, Table 5 show that using chain-of-thoughts mitigates this gap.
Different Input Lengths.
Figure 5(b) and Figure 5(e) show that the model can still do ICL for the examples with length equal to $4$. However, compared with the performance on examples with length equal to $3$ (Figure 5(c) and Figure 5(f)), the performance is worse. This may be because solving these length-4 examples requires more reasoning steps.
With/Without Rewriting the First Step.
Figure 8 shows that models trained with proofs that are rewritten has similar performance as models trained with the proofs that were rewritten (Figure 5). This suggests that rewriting the first step in the proof is not necessary for the model to acquire the ICL ability.
Model Size.
Figure 9 show that deeper models have better ICL performance. It aligns with the real-world observation that scaling helps model performance.
                                      $x_{1}$$x_{2}$$x_{4}$ $x_{3}$$r_{1}$ $x_{1}$$x_{3}$$x_{6}$ $x_{7}$ $\vdots\vdots$
(a) The proof tree a paragraph in the training dataset corresponds .
                       $x_{1}$$x_{2}$$x_{5}$ $x_{3}$$x_{4}$$x_{6}$ $r_{1}$
(b) A balanced tree for a downstream task sample.
Figure 4: Proof trees examples.
F.4 Hyper-parameters
We train our model using batch size 256, warm up ratio 5%, and we truncate the sequence length to 512 tokens and the default parameters for the optimizer. We use the implementation of GPT-2 by Hugging Face transformers v4.27.2. All models can be trained with 4 RTX 2080ti within 8 hours.
(a) balanced, length = 44, $y_{1},y_{2}$
(b) branching, length = 44, $y_{1},y_{2}$
(c) branching, length = 33, $y_{1},y_{2}$
(d) balanced, length = 33, $y_{3},y_{4}$
(e) branching, length = 44, $y_{3},y_{4}$
(f) branching, length = 33, $y_{3},y_{4}$
Figure 5: In-context learning accuracy with Calcutec when using different verbalizers ($y_{1},y_{2}$ or $y_{3},y_{4}$) and input lengths (3 or 4). The dotted lines represent the performance of unseen combinations described in §7.1.2, while the different colors represent the number of formulas each class ($v_{+}$ or $v_{-}$) is associated to. The main lines represent the average accuracy of 5 tasks. We plot the performance of each task in lighter colors.
(a) balanced, length = 44, $r_{1},r_{2}$
(b) branching, length = 44, $r_{1},r_{2}$
(c) branching, length = 33, $r_{1},r_{2}$
(d) balanced, length = 44, $r_{3},r_{4}$
(e) branching, length = 44, $r_{3},r_{4}$
(f) branching, length = 33, $r_{3},r_{4}$
Figure 6: In-context learning accuracy with Calcutec when no steps are dropped ($p_{skip}=0$).
(a) balanced, length = 44, $r_{1},r_{2}$
(b) branching, length = 44, $r_{1},r_{2}$
(c) branching, length = 33, $r_{1},r_{2}$
(d) balanced, length = 44, $r_{3},r_{4}$
(e) branching, length = 44, $r_{3},r_{4}$
(f) branching, length = 33, $r_{3},r_{4}$
Figure 7: In-context learning accuracy with Calcutec without rewriting the first step to include contain the premise of the proof.
(a) balanced, $r_{1},r_{2}$, 25000 examples
(b) balanced, $r_{1},r_{2}$, 10000 examples
(c) balanced, $r_{1},r_{2}$, 5000 examples
(d) balanced, $r_{3},r_{4}$, 25000 examples
(e) balanced, $r_{3},r_{4}$, 10000 examples
(f) balanced, $r_{3},r_{4}$, 5000 examples
(g) branching, $r_{1},r_{2}$, 25000 examples
(h) branching, $r_{1},r_{2}$, 10000 examples
(i) branching, $r_{1},r_{2}$, 5000 examples
(j) branching, $r_{3},r_{4}$, 25000 examples
(k) branching, $r_{3},r_{4}$, 10000 examples
(l) branching, $r_{3},r_{4}$, 5000 examples
Figure 8: In-context learning accuracy with different sizes of
(a) balanced, length = 44, 3 layers
(b) balanced, length = 44, 4 layers
(c) balanced, length = 44, 5 layers
(d) branching, length = 44, 3 layers
(e) branching, length = 44, 4 layers
(f) branching, length = 44, 5 layers
Figure 9: The in-context learning performance when using models with different model depths.
Branching Balanced
$r_{1},r_{2}$ $r_{3},r_{4}$ $r_{1},r_{2}$ $r_{3},r_{4}$
Task ICL CoT ICL CoT ICL CoT ICL CoT
Single 57.1 91.7 55.6 92.0 68.5 89.8 64.9 90.3
Double 53.5 76.3 51.1 77.1 58.5 76.1 56.2 75.8
Triple 53.0 73.0 51.7 73.4 57.0 68.2 54.2 67.0
Table 5: The 4-shot accuracy of in-context learning (ICL) versus chain-of-thoughts (CoT).
branching balance
$r_{1},r_{2}$ $r_{3},r_{4}$ $r_{1},r_{2}$ $r_{3},r_{4}$
#-shot 2 4 6 2 4 6 2 4 6 2 4 6
single 49.1 89.5 84.0 59.5 92.0 86.9 58.5 86.2 85.5 50.3 90.3 89.9
double 47.8 71.4 75.6 53.1 77.1 86.1 49.1 70.4 69.0 50.5 75.8 79.4
triple 46.7 65.7 70.7 50.6 73.4 79.4 46.0 60.2 61.4 49.8 67.0 70.4
Table 6: The CoT performance with 2, 4, or 6 examples.
Sample $N=\{n_{1},n_{2},n_{3},n_{4},n_{5}\}$ from either $\{0,1,\cdots,9\}$ or $\{6,7,\cdots,15\}$.
Sample four operators $O=\{o_{1},o_{2},o_{3},o_{4}\}$ from $\{+,\times\}$.
Initialize an empty list $S$ for the storing the steps.
while $\|N\|>1$ do
  Randomly sample an index $i$ of an operator from $O$. If $\times\in O$, then make sure that the sampled index corresponds to $\times$.
  Create a step $s\leftarrow[n_{i},n_{i+1},n_{i}o_{i}n_{i+1}\mod 16]$ and append $s$ to $S$.
  Remove $n_{i}$ and $n_{i+1}$ from $N$.
  Remove $o_{i}$ from $O$.
  Insert $n_{i}o_{i}n_{i+1}\mod 16$ into NN at position ii.
end while
return $S$.
Algorithm 2 Pseudo code for the generation process of a training example for the arithmetic task.
Given a list of steps $S=\langle s_{1},s_{2},\cdots,s_{4}\rangle$. (Each step is represented with a list, where the last number in the list is the right-hand side of an equation.)
Initialize n list $S^{\prime}=\langle s_{1}\rangle$ for the storing the randomly merged steps.
for $i\in\{2,3,4,5\}$ do
  Uniformly sample a number $r$ from $\{0,1\}$.
  if $r=0$ then
   Append $s_{i}$ to $S^{\prime}$.
  else
   if $S^{\prime}[-1][-1]\in s_{i}[:-1]$ then
     Remove $S^{\prime}[-1][-1]$ from $s_{i}$.
     $S^{\prime}[-1]\leftarrow\text{concatenate}(S^{\prime}[-1][:-1],s_{i})$
   end if
  end if
end for
return $S^{\prime}$.
Algorithm 3 Pseudo code for the generation process of a training example for the arithmetic task.
Appendix G Details of the Arithmetic Task
G.1 Training Data
When generating a training example, we first sample five numbers. These five numbers are either from $\{0,\cdots,9\}$ or $\{6,\cdots,15\}$. We then repeat the following process until there is only one number remaining in the list: We randomly pick two neighbor numbers from the list of five numbers and sum them and insert the result back into the list. This process results in a sequence of four steps, each of which is an equation that adds two numbers and produces one number. We outline the algorithm in Algorithm 2.
After we generate the four steps, we randomly merge steps into one step. We iterate through the steps. If the left-hand side current step contains the result of a previous step, we merge these two step into a single step at probability 0.5. For example, if the two steps are “1 + 2 = 3 . 3 + 4 = 7.” We merge these two steps as “1 + 2 + 4 = 7.” We outline the algorithm in Algorithm 3.
G.2 Hyper-parameters
We generate 80000 examples for training, 10000 for validation and 10000 for evaluation in the seen setup. For the unseen setup, we generate $6^{5}=7776$ examples. We train five Transformer models of the GPT-2-small architecture with five random seeds for 15 epochs. We use batch size 64 and the default training hyper parameters in HuggingFace transformers-4.46.3, i.e., with learning rate 5e-5, warm-up ratio 0, AdamW optimizer.
Appendix H Dataset License
  • SST-2: MIT
  • MR: unavailable
  • CR: unavailable
  • Subj: unavailable